OpenStudy (anonymous):

need help with system: -14=-2=y-7x 2x=10y +4 PLEASE HELP!!!!

OpenStudy (anonymous):

What do you need to do with the system?

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

Is the system\[\begin{matrix}-14 - 2 = y-7x \\ 2x = 10y + 4\end{matrix}\]

OpenStudy (anonymous):

Do you have to use any particular method?

OpenStudy (anonymous):

sorry.... it is: -14=-20y-7x 10y +4 = 2x

OpenStudy (anonymous):

solve by elimination....

OpenStudy (anonymous):

(sigh...)

OpenStudy (anonymous):

Weird. I replied to this but my answer isn't showing up. Let me re-type it real fast.

OpenStudy (anonymous):

'k

OpenStudy (anonymous):

First, we need to solve the first equation for x. We get\[x = -(20/7)y + 2\] Then we substitute this expression into the second equation. This yields \[10y - 2(-(20/7)y + 2) = -4\]Solve this for y and plug the value of y back into our expression for x. Let me know what you get.

OpenStudy (anonymous):

x=34 y=?